# Robust image classification with multi-modal large language models

### v1

@@ IV-A Experimental Setup
Datasets. We evaluate Multi-Shield on two widely-used benchmark datasets: CIFAR-10 [32] and ImageNet [33]. For CIFAR-10, we use the entire test set, which consists of $10,000$ images, and conduct experiments using a batch size of 32. For ImageNet, we randomly select a subset of $1,000$ images from the validation set, attacking batches of 8 samples at a time.
@@
Classifiers. For the image classifier, we source baseline and top state-of-the-art $\ell_{\infty}$ robust models from the RobustBench library  [34]. For CIFAR-10, we evaluate six models denoted with C1-C6. C1 [34] is a non-robust model. C2 [35], C4 [36] and C5 [37] employ adversarial training combined with data augmentation. C3 [38] utilizes a robust ensemble of models, while C6 [39] reformulates adversarial training to maximize the margin for more vulnerable samples, enhancing overall robustness. For ImageNet, we select four robust models, denoted with I1-I4, incorporating adversarial training. I1 [40] and I4 [41] utilize a ViT architecture, I2 [42] employs a Swin Transformer, and I3 [41] uses the ConvNeXt architecture.
@@
Multishield Construction. The Multi-Shield module combines a CLIP model with a standard image classifier. For CIFAR-10, we use a ViT-B visual encoder fine-tuned for CIFAR-10 classification [43], and for ImageNet, a pre-trained ViT-L [44], both paired with their respective text encoders from Hugging Face [45]
@@
Attack. We conduct the experiments using AutoAttack [7], a state of the art algorithm. We use both a naive (non-adaptive) version of AutoAttack and an adaptive variant, which takes into account Multi-Shield ’s rejection mechanism. Lastly, we set AutoAttack’s perturbation size to $\varepsilon=8/255$, as typical reference for robustness evaluation [34]
@@
Evaluation Metrics. In our experiments, we assess Multi-Shield’s performance using several key metrics. Clean Accuracy measures the accuracy on unperturbed inputs, providing a baseline for the model’s performance on pristine data. In this case, rejected samples are considered errors. The difference between the Clean Accuracy without any defense in place and Multi-Shield’s Clean Accuracy corresponds to the fraction of wrongly rejected samples (i.e., false positives). Robust Accuracy evaluates how well Multi-Shield performs on adversarial inputs by accounting for correct predictions and cases where it abstains from making a prediction. In fact, the desired model behavior in the presence of adversarial examples is to avoid misclassification. The Rejection Ratio represents the percentage of inputs for which Multi-Shield opts to abstain, reflecting the effectiveness of the rejection mechanism in identifying adversarial examples. The difference between the Robust Accuracy and Rejection Ratio corresponds to the fraction of correctly classified test inputs. Finally, we also measure the Execution Time, which indicates the computational cost of performing on Multi-Shield both the naive and adaptive version of the attack. All experiments have been run on a NVIDIA A100 Tensor Core GPU and two Intel® Xeo® Gold 6238R processors.
@@ IV-B Experimental Results
We present in Table I the results on the effectiveness of Multi-Shield in detecting adversarial examples across three distinct attack scenarios: (1) the baseline image classifiers is the target of attack without Multi-Shield detection (Baseline); (2) Multi-Shield is integrated during the final prediction while the image classifier is under attack (Multi-Shield); and (3) an adaptive attacker targets both the image classifier and Multi-Shield, providing a worst-case evaluation of the defense mechanism (Multi-Shield - Adaptive Attack). Finally, we integrate our effectiveness analysis with Figure 2 showing the performance of Multi-Shield against adversaries of varying strengths $\varepsilon$.
(a) C2 [35]
(b) C6 [39]
(c) I2 [42]
Figure 2: Curves showing the baseline’s robust accuracy alongside Multi-Shield’s robust accuracy and rejection ratio under attacks with varying perturbation sizes ($\varepsilon$) across different robust models.
@@
Multi-Shield Effectiveness. We analyze the robustness of both robust and non-robust pre-trained models without Multi-Shield (Baseline) and assess the impact of integrating Multi-Shield to enable rejection of anomalous samples (Multi-Shield). The results in Table I confirm that Multi-Shield significantly enhances robust accuracy across all models. For example, we observe that Multi-Shield raises the robust accuracy of the non-robust model C1 from $0\%$ (Baseline) to $91.8\%$ (Multi-Shield), successfully detecting nearly all adversarial examples and abstaining from incorrect predictions. The robustness boost with Multi-Shield is notable even on robust models, with an average improvement of $32\%$ on CIFAR-10 and $65\%$ on ImageNet (Multi-Shield). These results highlight Multi-Shield’s effectiveness as an additional security layer to increase model resilience without requiring retraining of the underlying image classifier. The only trade-off is a slight reduction in clean accuracy, averaging $1.8\%$ on CIFAR-10 and $6\%$ on ImageNet.
@@
Impact of Attack Strength on Multi-Shield. We evaluate Multi-Shield against attackers of varying strengths (i.e., at different values of $\varepsilon$), demonstrating how robust accuracy and rejection ratio evolve as attack power increases. These trends are illustrated across three robust models (C2,C6,I2) in Figure 2. Firstly, in all three plots, the gap between the robust accuracy of Multi-Shield (solid blue line) and the baseline (dashed gray line) widens as $\varepsilon$ increases, highlighting Multi-Shield’s superior robustness in response to increasing attack strength. Secondly, the rejection ratio (red line) steadily increases in all three models, showcasing Multi-Shield’s ability to detect and reject stronger adversarial examples that mislead the baseline. These results become more evident in the ImageNet model (2(c)), where robust accuracy remains relatively more resilient while rejections rise significantly.
@@
Adaptive Attacker. We conclude our analysis by evaluating Multi-Shield’s performance under the worst-case scenario of an adaptive attacker who knows everything about the target model and exploits knowledge of the defense mechanism to bypass it [31]. Specifically, we assume an adaptive attacker with complete knowledge of the rejection mechanism, including Multi-Shield’s image classifier, rejection mechanism, and inner workings (i.e., parameters and class prompts). Firstly, looking at Table I rightmost scenario (Multi-Shield - Adaptive Attack), Multi-Shield significantly improves robust accuracy of the non-robust C1 by 20%. For robust models, Multi-Shield enhances robustness by an average of $3.3\%$ on CIFAR-10 and $5\%$ on ImageNet. This indicates that attacking Multi-Shield, even under worst-case evaluation, requires significantly more effort and introduces further difficulties. Specifically, the optimization process for generating adaptive adversarial examples becomes more difficult because the constraint in Eq. (missing) 11 demands the attack to simultaneously mislead both the unimodal image classifier and the multi-modal CLIP classifier in Multi-Shield. Finally, adaptive attacks contribute to a complete system evaluation. However, the complete knowledge demand of the scenario and harder-to-optimize objective function still pose challenges for the attacker.
V Conclusion
@@
In this work, we introduce Multi-Shield, a novel defense that implements a rejection mechanism based on two core principles: (i) integration of multi-modal information and (ii) interaction between adversarial training and a multi-modal adversarial detector. The first principle is realized by analyzing visual features and their semantic alignment with the textual prompts for the classes. In this way, Multi-Shield enables the identification and rejection of adversarial examples when classification presents uncertainties or semantic inconsistencies. It leverages agreement in predictions between an image classifier and a multi-modal model to decide whether to abstain from making a prediction. Secondly, Multi-Shield complements existing defenses that rely on adversarial training, providing an additional security layer against adversarial attacks. Our experiments show that Multi-Shield consistently detects adversarial examples that mislead image classifiers and remains effective even under worst-case adaptive attacks. For future work, we aim to explore alternative methods for crafting more descriptive class label prompts, investigate their impact on robustness, and extend adversarial training to the multi-modal model to further strengthen the defense.
Acknowledgment
@@
This work has been partially supported by the EU—NGEU National Sustainable Mobility Center (CN00000023), Italian Ministry of University and Research Decree n. 1033—17/06/2022 (Spoke 10), and project SERICS (PE00000014) under the MUR National Recovery and Resilience Plan funded by the European Union—NextGenerationEU. Lastly, this work was carried out while Dario Lazzaro was enrolled in the Italian National Doctorate on Artificial Intelligence run by the Sapienza University of Rome in collaboration with the University of Genoa.

### v2

@@ IV-A Experimental Setup
Datasets. We consider six datasets: CIFAR-10 [Krizhevsky, 2009], ImageNet [Krizhevsky et al., 2012], Caltech-101 [Fei-Fei et al., 2004], Food-101 [Bossard et al., 2014], Oxford-IIIT Pets Parkhi et al. [2012], STL-10 Coates et al. [2011]. For CIFAR-10, we use all 10,000 test images. For ImageNet, we randomly sample $5000$ validation images. We evaluate the full validation sets for Caltech-101, Oxford-IIIT Pets, and STL-10, and sample $10000$ images for Food-101.
@@
Classifiers. For the image classifier, we source baseline and top state-of-the-art $\ell_{\infty}$ robust models from the RobustBench library [Croce et al., 2021]. For CIFAR-10, we evaluate six models denoted with C1-C6. C1 [Croce et al., 2021] is a non-robust model. C2 [Carmon et al., 2019],C4 [Gowal et al., 2021] and C5 [Addepalli et al., 2022] employ adversarial training combined with data augmentation. C3 [Chen et al., 2020] utilizes a robust ensemble of models, while C6 [Xu et al., 2023] reformulates adversarial training to maximize the margin for more vulnerable samples, enhancing overall robustness. For ImageNet, we select four robust models, denoted with I1-I4, incorporating adversarial training. I1 [Debenedetti et al., 2023] and I4 [Singh et al., 2024] utilize a ViT architecture, I2 [Liu et al., 2024] employs a Swin Transformer, and I3 [Singh et al., 2024] uses the ConvNeXt architecture. For experiments on Caltech-101, Oxford-IIIT Pets, STL-10, and Food-101, we fine-tune three ImageNet-pretrained models—ResNet-50 (M1), WideResNet50-2 (M2), and ViT-B/16 (M3). For each model, we replace the final classification layer and fine-tune the entire model for each dataset over $50$ epochs using stochastic gradient descent with a learning rate of 0.001 and momentum of $0.9$.
@@
Multishield Construction. The Multi-Shield module combines a CLIP model with a standard image classifier. For CIFAR-10, we use a ViT-B visual encoder fine-tuned for CIFAR-10 classification [Tang et al., 2024], and for ImageNet, a pre-trained ViT-L [Li et al., 2023], both paired with their respective text encoders from Hugging Face Wolf et al. [2019]
@@
Attack. We conduct the experiments using AutoAttack [Croce and Hein, 2020], a state of the art algorithm. We use both a naive (non-adaptive) version of AutoAttack and an adaptive variant, which takes into account Multi-Shield ’s rejection mechanism. Lastly, we set AutoAttack’s perturbation size to $\varepsilon=8/255$, as typical reference for robustness evaluation [Croce et al., 2021]
@@
Evaluation Metrics. In our experiments, we assess Multi-Shield’s performance using several key metrics. Clean Accuracy measures the accuracy on unperturbed inputs, providing a baseline for the model’s performance on pristine data. In this case, rejected samples are considered errors. The difference between the Clean Accuracy without any defense in place and Multi-Shield’s Clean Accuracy corresponds to the fraction of wrongly rejected samples (i.e., false positives). Robust Accuracy evaluates how well Multi-Shield performs on adversarial inputs by accounting for correct predictions and cases where it abstains from making a prediction. In fact, the desired model behavior in the presence of adversarial examples is to avoid misclassification. The Rejection Ratio represents the percentage of inputs for which Multi-Shield opts to abstain, reflecting the effectiveness of the rejection mechanism in identifying adversarial examples. The difference between the Robust Accuracy and Rejection Ratio corresponds to the fraction of correctly classified test inputs. Lastly, all experiments have been run on a NVIDIA A100 Tensor Core GPU with 40 GB of VRAM.
@@ IV-B Experimental Results
We present in Table I the results on the effectiveness of Multi-Shield in detecting adversarial examples across three distinct attack scenarios: (1) the baseline image classifiers is the target of attack without Multi-Shield detection (Baseline); (2) Multi-Shield is integrated during the final prediction while the image classifier is under attack (Multi-Shield); and (3) an adaptive attacker targets both the image classifier and Multi-Shield, providing a worst-case evaluation of the defense mechanism (Multi-Shield - Adaptive Attack). We then integrate our analysis with Figure 2 showing the performance of Multi-Shield against adversaries of varying strengths $\varepsilon$. Lastly, we compare Multi-Shield with two state-of-the-art rejection defenses (i.e., Melis et al. [2017] and Sotgiu et al. [2020]) in Table II.
TABLE II: Comparison of Multi-Shield, NR Melis et al. [2017], and DNR Sotgiu et al. [2020] over five runs on 1,000-sample subsets from CIFAR-10. The rejection thresholds for NR and DNR are set to roughly match Multi-Shield’s accuracy. Robust Acc. (A) refers to accuracy under an adaptive attack.
Defense Model Clean acc. Robust acc. Robust acc. (A) Model Clean acc. Robust acc. Robust acc. (A)
NR C1 91.7 $\pm$ 0.002 3.2 $\pm$ 0.001 0.00 $\pm$ 0.00 C2 88.4 $\pm$ 0.01 75.6 $\pm$ 0.03 59.2 $\pm$ 0.04
DNR 91.5 $\pm$ 0.001 96.1 $\pm$ 0.004 0.00 $\pm$ 0.00 88.2 $\pm$ 0.006 76.3 $\pm$ 0.02 60.6 $\pm$ 0.05
Ours 92.7 $\pm$ 0.002 91.9 $\pm$ 0.006 20.2 $\pm$ 0.003 88.3 $\pm$ 0.007 91.7 $\pm$ 0.009 64.2 $\pm$ 0.05
NR C5 84.3 $\pm$ 0.02 72.1 $\pm$ 0.02 49.3 $\pm$ 0.05 C6 90.7 $\pm$ 0.007 75.9 $\pm$ 0.03 66.3 $\pm$ 0.04
DNR 84.2 $\pm$ 0.009 73.1 $\pm$ 0.04 50.8 $\pm$ 0.04 90.8 $\pm$ 0.009 76.0 $\pm$ 0.03 66.6 $\pm$ 0.03
Ours 83.8 $\pm$ 0.004 92.5 $\pm$ 0.008 58.9 $\pm$ 0.05 91.3 $\pm$ 0.002 88.8 $\pm$ 0.009 66.9 $\pm$ 0.07
NR C3 82.6 $\pm$ 0.003 66.1 $\pm$ 0.01 43.8 $\pm$ 0.03 C4 82.3 $\pm$ 0.006 78.7 $\pm$ 0.03 62.1 $\pm$ 0.05
Ours 84.5 $\pm$ 0.003 91.9 $\pm$ 0.02 55.0 $\pm$ 0.03 87.0 $\pm$ 0.002 93.3 $\pm$ 0.008 69.6 $\pm$ 0.06
@@
Multi-Shield Effectiveness. We analyze the robustness of pre-trained models without Multi-Shield (Baseline) and assess the impact of integrating Multi-Shield to enable rejection of anomalous samples (Multi-Shield). The results in Table I confirm that Multi-Shield serves as an effective additional security layer significantly enhancing robust accuracy across all models without requiring their retraining. For example, Multi-Shield raises the robust accuracy of non-robust models (i.e., C1, M1, M2, M3) from near $0\%$ (Baseline) to over $90\%$ across multiple datasets, effectively rejecting most adversarial examples. The robustness boost with Multi-Shield is notable even on robust models, with an average improvement of $32\%$ on CIFAR-10 and $65\%$ on ImageNet (Multi-Shield). The only trade-off is a slight reduction in clean accuracy, averaging approximately $1.8\%$ on CIFAR-10, $6\%$ on ImageNet and Caltech-101, $3\%$ on Food-101 and Oxford-IIIT Pets, and under $1\%$ on STL-10.
@@
Impact of Attack Strength on Multi-Shield. We evaluate Multi-Shield against attackers of varying strengths (i.e., at different values of $\varepsilon$), demonstrating how robust accuracy and rejection ratio evolve as attack power increases. These trends are illustrated in Figure 2 across four robust models (C2,C6,I2,I4) on CIFAR-10 (top two plots) and ImageNet (bottom two plots). Firstly, in all three plots, the gap between the robust accuracy of Multi-Shield (solid blue line) and the baseline (dashed gray line) widens as $\varepsilon$ increases, highlighting Multi-Shield’s superior robustness in response to increasing attack strength. Secondly, the rejection ratio (red line) steadily increases in all three models, showcasing Multi-Shield’s ability to detect and reject stronger adversarial examples that mislead the baseline. These results become more evident in the ImageNet models (2(c)-2(d)), where robust accuracy remains relatively more resilient while rejections rise significantly.
(a) C2 [Carmon et al., 2019]
(b) C6 [Xu et al., 2023]
(c) I2 [Liu et al., 2024]
(d) I4 [Singh et al., 2024]
Figure 2: Curves showing the baseline robust accuracy alongside Multi-Shield’s robust accuracy and rejection ratio under attacks with varying perturbation sizes ($\varepsilon$) across different robust models on CIFAR-10 (top two plots) and ImageNet (bottom two plots).
@@
Adaptive Attacker. We conclude our analysis by evaluating Multi-Shield under the worst-case scenario of an adaptive attacker who knows everything about the target model and exploits knowledge of the defense mechanism to bypass it [Carlini et al., 2019]. Specifically, we give the strong assumption of an adaptive attacker with complete knowledge of the rejection mechanism, including Multi-Shield’s image classifier, rejection mechanism, and inner workings (i.e., parameters and class prompts) to have a complete system evaluation. Firstly, looking at the rightmost scenario in Table I (Multi-Shield - Adaptive Attack), Multi-Shield provides substantial improvements in robust accuracy across all non-robust models (C1, M1, M2, M3). While the gains are already significant for CIFAR-10 ($20.4\%$) and Oxford-IIIT Pets ($22.1\%$), they become even larger for Caltech-101 ($59.1\%$), Food-101 ($55.9\%$), and STL-10 ($43\%$). For robust models, Multi-Shield enhances robustness by an average of $3.3\%$ on CIFAR-10 and $5\%$ on ImageNet. This indicates that attacking Multi-Shield, even under worst-case evaluation, requires significantly more effort and introduces further difficulties for the attacker. The constraint in Eq. (missing) 11 forces the attack to simultaneously deceive both the unimodal image classifier and the multimodal CLIP classifier, complicating the optimization process. We conclude that Multi-Shield serves as an additional layer of defense to enhances the robustness of all considered models, both under non-adaptive and adaptive (worst-case) attack settings, across all datasets.
Comparison with other Rejection Defenses. We compare our method with state-of-the-art rejection defenses, including Neural Rejection (NR) [Melis et al., 2017] and Deep Neural Rejection [Sotgiu et al., 2020], following the experimental setups outlined in their original works. Experiments are conducted on five distinct subsets of $1000$ CIFAR-10 samples, reporting the average and standard deviation of results. The results, reported in Table II, showcase that Multi-Shield consistently outperforms NR and DNR in robustness across various models. Notably, while DNR achieves high robustness under C1 (e.g., $96.1\%$), it fails entirely under adaptive attacks, dropping to $0.00\%$. In contrast, Multi-Shield maintains significantly higher robustness against adaptive attackers (e.g., $20.2\%$ vs. $0.00\%$ for C1). Similar improvements are observed across other robust models, and the low standard deviation across multiple runs confirms the consistency of our results, reinforcing the validity of our conclusions. On C5, Multi-Shield achieves $58.9\%$ robust accuracy under adaptive attacks, outperforming NR and DNR by approximately $8$ and $9$ percentage points, respectively. Likewise, on C4, Multi-Shield reaches $69.6\%$ robust accuracy, surpassing NR ($62.1\%$) by $7.5\%$. Beyond robustness, Multi-Shield is also computationally less demanding. Both NR and DNR require training one or multiple one-vs-all SVMs for each class, leading to substantial memory and computational overhead. DNR, in particular, necessitates extreme batch size reductions (e.g., batch size = 1) to fit within 40GB of VRAM when applied to CIFAR-10, making it impractical for larger datasets (e.g., ImageNet) or more complex models (e.g., C3, C4). We thus conclude that these comparisons further emphasize the benefits of Multi-Shield as an additional security layer for robust classification systems, enhancing resilience against adversarial threats while maintaining high accuracy and computational efficiency.
V Conclusions, Limitations, and Future Works
@@
In this work, we introduce Multi-Shield, a novel defense that implements a rejection mechanism based on two core principles: (i) integration of multimodal information and (ii) interaction between adversarial training and a multimodal adversarial detector. The first principle is realized by analyzing visual features and their semantic alignment with the textual prompts for the classes. In this way, Multi-Shield enables the identification and rejection of adversarial examples when classification presents uncertainties or semantic inconsistencies. It leverages agreement in predictions between an image classifier and a multimodal model to decide whether to abstain from making a prediction. Secondly, Multi-Shield complements existing defenses that rely on adversarial training, providing an additional security layer against adversarial attacks. Our experiments show that Multi-Shield consistently detects adversarial examples that mislead image classifiers and remains effective even under worst-case adaptive attacks. Despite its strengths, Multi-Shield may struggle to detect adversarial inconsistencies in datasets with abstract or meaningless labels, such as “Class 1” or “Class 2,” as these provide no meaningful semantic cues for the multimodal model to align textual and visual data. For future work, we aim to explore alternative methods for crafting more descriptive class label prompts, investigate their impact on robustness, and extend adversarial training to the multimodal model to further strengthen the defense.
@@
Acknowledgments. This work has been partially supported by the EU—NGEU National Sustainable Mobility Center (CN00000023), Italian Ministry of University and Research Decree n. 1033—17/06/2022 (Spoke 10), and project SERICS (PE00000014) under the MUR National Recovery and Resilience Plan funded by the European Union—NextGenerationEU. Lastly, this work was carried out while Dario Lazzaro was enrolled in the Italian National Doctorate on Artificial Intelligence run by the Sapienza University of Rome in collaboration with the University of Genoa.